# All That Glitters is Not Novel: Plagiarism in AI Generated Research

### v1

@@ Abstract
Automating scientific research is considered the final frontier of science. Recently, several papers claim autonomous research agents can generate novel research ideas. Amidst the prevailing optimism, we document a critical concern: a considerable fraction of such research documents are smartly plagiarized. Unlike past efforts where experts evaluate the novelty and feasibility of research ideas, we request $13$ experts to operate under a different situational logic: to identify similarities between LLM-generated research documents and existing work. Concerningly, the experts identify $24\%$ of the $50$ evaluated research documents to be either paraphrased (with one-to-one methodological mapping), or significantly borrowed from existing work. These reported instances are cross-verified by authors of the source papers. Problematically, these LLM-generated research documents do not acknowledge original sources, and bypass inbuilt plagiarism detectors. Lastly, through controlled experiments we show that automated plagiarism detectors are inadequate at catching deliberately plagiarized ideas from an LLM. We recommend a careful assessment of LLM-generated research, and discuss the implications of our findings on research and academic publishing. [^1] Our code, along with expert-provided scores and explanations for each proposal, is available at: https://anonymous.4open.science/r/AI-Papers-Plagiarism-ECCA
@@ 1 Introduction
Our expert-led analysis reveals that $14.0\%$ of LLM-generated research documents receive a score of $5$ using the rubric in Table 1, and another $10.0\%$ documents receive a score of $4$. We consider both scores $4$ and $5$ as instances of plagiarism, totaling $24.0\%$ of proposals with noticeable plagiarism. These scores are verified by emailing the authors of referenced papers. When including claims where source paper authors are unreachable, these numbers increase: $18.0\%$ of proposals receive a score of $5$ and $18.0\%$ receive a score of $4$, amounting to $36.0\%$ of the examined proposals. Notably, several previously showcased exemplars of LLM generated research (Si et al., 2024; Lu et al., 2024a) are found to be plagiarized. Our experimental setup and key findings are illustrated in Figure 1.
@@ 3 Background: Generating Proposals
The research proposal generation process consists of six sequential steps, with Claude $3.5$ Sonnet (Anthropic, 2024) being the backbone LLM. For a given topic, the first step uses a retrieval-augmented generation (RAG) system to retrieve and rank relevant papers using the Semantic Scholar API. The second step generates initial seed ideas using these retrieved papers, while the third step involves deduplication using text embeddings, retaining about $5\%$ of the original ideas. The fourth step expands these seed ideas into detailed project proposals, and the fifth step implements a Swiss tournament system where proposals compete in pairwise comparisons over five rounds to identify the strongest candidates. We refer readers to Si et al. (2024) for additional details on these steps.
The final step, most relevant to our work, attempts to detect potential plagiarism through Semantic Scholar Augmented Generation (SSAG). First, an LLM iteratively generates queries for the Semantic Scholar API to find papers similar to the proposal’s content, using results from previous iterations as context to inform each new query. This search process continues either until they collect $100$ papers or reach $10$ iterations, whichever comes first. The retrieved papers are then scored by an LLM based on their relevance, narrowing down to $10$ most similar papers. Further, Claude $3.5$ Sonnet performs pairwise comparisons between the proposal and each of these top $10$ papers. The LLM is prompted to determine if the research proposal and the retrieved paper are substantially similar, discarding the proposal if they are. This process removes only about $1\%$ of the generated proposals.
@@
For our study’s implementation, Si et al. (2024)’s method is slightly modified to optimize computational costs while maintaining effectiveness. Instead of generating $4000$ ideas per research topic, we generate $500$ ideas per topic. This reduces our initial pool from around $200$ unique ideas per topic (as in the original paper (Si et al., 2024)) to $138$, resulting in significant computational cost savings while only reducing unique ideas by $31\%$. All other components of the pipeline, including the plagiarism detection and ranking systems, are exactly the same as Si et al. (2024)’s original implementation.
The choice to use this specific methodology (Si et al., 2024) for generating research proposals is motivated by several factors. First, it is representative of fundamental prompt engineering techniques used across various research idea generation systems (Li et al., 2024a; Lu et al., 2024a; Baek et al., 2024; Li et al., 2024c; Wang et al., 2023; Yang et al., 2023; Weng et al., 2024). Second, as discussed in §2, this study conducts the most comprehensive evaluation to date, and finds that human participants judge LLM-generated research more novel than those by humans. Third, this method uses minimal prompt engineering. Other approaches like Li et al. (2024a) use sophisticated prompt engineering that function as “natural language programs,” potentially incorporating human creativity. By keeping prompts simple, we can better assess the raw capabilities of LLMs themselves.
@@ 4.1 Expert Evaluation Design
Our experimental setup involves generating five research proposals for each of the twelve topics listed in Table 4. These NLP topics are determined by asking participants to describe their areas of expertise, ensuring evaluators have in-depth familiarity with the current literature for detecting potential plagiarism. Each participant evaluates three out of five proposals that aligns with their domain knowledge, resulting in $3\times 12=36$ total proposals.
In addition to these $36$ proposals, we include fourteen previously-showcased exemplars—four research proposals from Si et al. (2024) and ten research papers from Lu et al. (2024a), bringing the total to $50$ research documents. These fourteen exemplars were evaluated by two experts. We use convenience sampling to recruit participants who are actively conducting research in NLP through our professional networks. Our participant pool comprises experts from $5$ universities and $2$ industrial labs, with $69\%$ being Ph.D. students or recent graduates and $31\%$ being associate researchers in industrial or academic labs. [^3] Our experts are affiliated with University of Washington, Carnegie Mellon University, Harvard University, University of Southern California, Indian Institute of Science, Allen Institute for Artificial Intelligence (Ai2) and Adobe, Inc. Unlike prior studies where participants assess LLM-generated research documents without suspecting plagiarism, our participants are instructed to actively search for potential overlaps with existing work.
@@ 5 Results
Our code and detailed findings, including all evaluated proposals and author verifications, are available at https://anonymous.4open.science/r/AI-Papers-Plagiarism-ECCA.
@@ 5.1 Expert Evaluation Results
Our expert evaluation reveals plagiarism in LLM-generated research documents, as shown in Table 2. Similarity scores across verified claims highlight substantial content misappropriation at various levels of severity. Since our evaluation is limited by our participants’ time constraints and the laborious manual effort required for thorough plagiarism checks, our results likely represent a lower bound on actual plagiarism rates. Interestingly, of the four exemplars presented in Si et al. (2024), one received a similarity score of $5$ and another received a score of $4$, while among the ten exemplars in Lu et al. (2024a), two received scores of $5$ and one received score $4$—all of these ratings are cross-verified by the source papers’ authors. It is crucial to note the documents that receive similarity scores of $5$ in our evaluation passed through SSAG plagiarism detection checks, indicating serious limitations of such systems (elaborated in §5.3).
@@
5.2 A Case Study
As shown in Figure 2, the proposed methodology exhibits a clear one-to-one mapping with the original paper. Each component of the LLM-generated proposal corresponds to specific sections in the source paper, albeit with skillfully reworded descriptions. The proposal proposes the same technical approach while using different terminology (e.g., “resonance graph” instead of “weighted adjacency matrix”) and restructuring the presentation. This may be interpreted as adversarial behavior, where the LLM has learned to disguise existing work as novel research through careful rewording. Notably, expert reviewers in Si et al. (2024)’s study do not identify this plagiarism, likely because their evaluation focuses on assessing novelty and feasibility rather than actively searching for potential sources of plagiarism. In contrast, our study’s participants, operating under different situational logic (Popper, 2013; Hoover, 2016) that presumes potential plagiarism, are able to identify the original paper. A similar analysis of plagiarism in a showcased paper from Lu et al. (2024a) is provided in Appendix A. This case-study, along with results of expert evaluation presented in §5.1, showcases the limitations of previous human evaluations of LLM generated research ideas—without a skeptical eye for plagiarism, experts may be fooled into considering LLM generated research ideas as novel.
5.3 Performance of Plagiarism Detectors
@@
The SSAG approach, a common plagiarism detection method in research document generation systems (Si et al., 2024; Lu et al., 2024a; Li et al., 2024a), attains moderate performance ($51.3\%$ for Claude, $68.5\%$ for GPT-4o). The considerable gap between SSAG performance and oracle access indicates that retrieving relevant papers, not determining similarity, is the bottleneck.
The inadequate performance of SSAG is particularly concerning for research agent systems that rely on it (Si et al., 2024; Lu et al., 2024a; Li et al., 2024a), especially given our test scenario represents a much easier challenge—our proposals are deliberately plagiarized from single papers, while research agent systems are specifically designed to produce allegedly novel content. (An illustrative example of successful plagiarism detection using the SSAG from our synthetic dataset is provided in Appendix B.)
@@ 6 Discussion
Citing Relevant Work Is Challenging.
Figure 3: PCA projection of title embeddings for human-written papers and LLM-generated proposals on the topic “Novel AI-assisted formal proof generation methods”. LLM-generated proposals occupy a more confined region, indicating less diversity in outputs.
Limited Diversity in Proposals.
Prior research notes limited diversity in LLM-generated research proposals (Si et al., 2024), which our analysis confirms. For each research topic in Table 4, we analyze titles of $100$ human-written papers and $100$ LLM-generated proposals. [^5] Due to deduplication and filtering in (Si et al., 2024), some topics had $<100$ proposals, resulting in $2370$ samples rather than the maximum of $2400$ ($200$ samples $\times$ $12$ topics). Using the all-MiniLM-L6-v2 sentence transformer (Reimers and Gurevych, 2020), we generate embeddings for these titles. Calculating cluster spread via mean squared distance from centroids, we find that despite aggressive deduplication strategies (Si et al., 2024), LLM-generated titles are more tightly clustered—the ratio of LLM to human cluster spreads averages $0.76$ across topics. Figure 3 illustrates this pattern, showing that LLM proposals occupy a more confined region in the embedding space. While this analysis is limited since titles only reflect broad research directions rather than specific ideas, the clear clustering pattern suggests current LLM systems explore a narrower range of research directions compared to humans, undermining their utility for scientific innovation.
Proposal Titles Are Easily Distinguishable.
Beyond being tightly clustered, LLM-generated proposals also appear to follow distinctive patterns in their research directions. In a preliminary analysis using $100$ human-written papers and $100$ LLM-generated proposals per topic from Table 4, a simple logistic regression model with bag-of-words features on a $60$:$40$ train-test split achieves $94.2\%$ accuracy in classifying titles.5 While this basic classification experiment has limitations as titles only capture broad research directions, it provides initial evidence that LLM-generated research follows predictable patterns. However, developing reliable detection methods will require significant further research and is beyond the scope of this work.
@@ 7 Limitations
Our expert evaluation design has some limitations. First, we ask our human participants to presume adversarial plagiarism and actively search for it. This may introduce confirmation bias, leading them to give higher similarity scores in order to complete their task. We indeed find $4$ instances where authors of source-papers over-turned the score downwards with difference of $2$ or more. However, our (author-verified) scores already account for these adjustments, and our findings remain independently verifiable through our open-sourced results. Second, we provide our human-participants with $5$ research proposals and ask them to choose any $3$. While this gives flexibility to our participants, since even in a single topic there can be sub-topics that participants may not be familiar with, it may introduce sampling bias.
@@
Title-only Analysis.
Our clustering and classification experiments (§6) rely solely on paper titles as proxies for broad research directions and topics, rather than using full text. While this approach may seem limited, we deliberately avoid using complete documents because LLM-generated proposals and human-written papers follow different organizational structures. These structural differences would confound any full-text analysis, making titles the most comparable unit for exploring patterns in research directions. These two experiments should be viewed as preliminary explorations rather than methodological recommendations.

### v2

@@ Abstract
Automating scientific research is considered the final frontier of science. Recently, several papers claim autonomous research agents can generate novel research ideas. Amidst the prevailing optimism, we document a critical concern: a considerable fraction of such research documents are smartly plagiarized. Unlike past efforts where experts evaluate the novelty and feasibility of research ideas, we request $13$ experts to operate under a different situational logic: to identify similarities between LLM-generated research documents and existing work. Concerningly, the experts identify $24\%$ of the $50$ evaluated research documents to be either paraphrased (with one-to-one methodological mapping), or significantly borrowed from existing work. These reported instances are cross-verified by authors of the source papers. Experts find an additional $32\%$ ideas to partially overlap with prior work, and a small fraction to be completely original. Problematically, these LLM-generated research documents do not acknowledge original sources, and bypass inbuilt plagiarism detectors. Lastly, through controlled experiments we show that automated plagiarism detectors are inadequate at catching plagiarized ideas from such systems. We recommend a careful assessment of LLM-generated research, and discuss the implications of our findings on academic publishing. [^1] Our code, along with expert-provided scores and explanations for each proposal, is available at: https://github.com/tarun360/AI-Papers-Plagiarism/
@@ 1 Introduction
Our expert-led analysis reveals that $14.0\%$ of LLM-generated research documents receive a score of $5$ and another $10.0\%$ documents receive a score of $4$. We consider both scores $4$ and $5$ as instances of plagiarism, totaling $24.0\%$ of proposals with noticeable plagiarism. These scores are verified by emailing the authors of referenced papers. When including claims where source paper authors are unreachable, these numbers increase: $18.0\%$ of proposals receive a score of $5$ and $18.0\%$ receive a score of $4$, amounting to $36.0\%$ of the examined proposals. Notably, several previously showcased exemplars of LLM generated research (Si et al., 2024; Lu et al., 2024a; Yamada et al., 2025) are either found to be plagiarized or substantially similar to existing work. Our setup and key findings are illustrated in Figure 1.
@@ 3 Background: Generating Proposals
The research proposal generation process consists of six sequential steps, with Claude $3.5$ Sonnet (Anthropic, 2024) being the backbone LLM. For a given topic, the first step uses a retrieval-augmented generation (RAG) system to retrieve and rank relevant papers using the Semantic Scholar API. The second step generates initial seed ideas using these retrieved papers, while the third step involves deduplication using text embeddings, retaining about $5\%$ of the original ideas. The fourth step expands these seed ideas into detailed project proposals containing a title, problem statement, motivation, proposed method, experiment plan, test cases, and a fall back plan. The fifth step implements a Swiss tournament system where proposals compete in pairwise comparisons over five rounds to identify the strongest candidates. We refer readers to Si et al. (2024) for additional details on these steps.
The final step, most relevant to our work, attempts to detect potential plagiarism through Semantic Scholar Augmented Generation (SSAG). First, an LLM iteratively generates queries for the Semantic Scholar API to find papers similar to the proposal’s content, using results from previous iterations as context to inform each new query. This search process continues either until they collect $100$ papers or reach $10$ iterations, whichever comes first. The retrieved papers are then scored by an LLM based on their relevance, narrowing down to $10$ most similar papers. Further, Claude $3.5$ Sonnet performs pairwise comparisons between the proposal and each of these top $10$ papers. The LLM determines if the proposal and the retrieved paper are substantially similar, discarding the proposal if they are. This process removes only about $1\%$ of the generated proposals (Si et al., 2024).
@@
For our study’s implementation, Si et al. (2024)’s method is slightly modified to optimize computational costs while maintaining effectiveness. Instead of generating $4000$ ideas per research topic, we generate $500$ ideas per topic. This reduces our initial pool from around $200$ unique ideas per topic (as in the original paper (Si et al., 2024)) to $138$, resulting in significant computational cost savings while only reducing unique ideas by $31\%$. This non-linear relationship between generated and unique ideas can be explained by the fact that as research agents generate more ideas, the percentage of unique ideas decreases and eventually plateaus, offering diminishing returns (Si et al., 2024). All other components of the pipeline, including the plagiarism detection and ranking, are exactly the same as Si et al. (2024)’s original implementation.
The choice to use this methodology (Si et al., 2024) for generating research proposals has several motivations. First, it is representative of prompt engineering techniques used across various research idea generation systems (Li et al., 2024a; Lu et al., 2024a; Baek et al., 2024; Li et al., 2024c; Wang et al., 2023; Yang et al., 2023; Weng et al., 2024). Second, as discussed in §2, this study conducts the most comprehensive evaluation to date, and finds that human participants judge LLM-generated research more novel than those by humans. Third, this method uses minimal prompt engineering. Other approaches like Li et al. (2024a) use sophisticated prompt engineering that function as “natural language programs,” potentially incorporating human creativity. By keeping prompts simple, we can better assess LLMs’ raw capabilities.
@@ 4.1 Expert Evaluation Design
Our experimental setup involves generating five research proposals for each of the twelve topics listed in Table 5. These NLP topics are determined by asking participants to describe their areas of expertise, ensuring evaluators have in-depth familiarity with the current literature for detecting potential plagiarism. Twelve participants each evaluate $3$ out of $5$ proposals that align with their domain knowledge, resulting in $3\times 12=36$ total proposals.
In addition to these $36$ proposals, we include fourteen previously-showcased exemplars—four research proposals from Si et al. (2024) and ten research papers from Lu et al. (2024a), bringing the total to $50$ research documents. Our study includes $13$ experts in total: $12$ experts evaluate the 36 proposals, and $2$ experts evaluate the $14$ exemplars, with one expert being common among both evaluation tasks. We use convenience sampling to recruit participants who are actively conducting research in NLP through our professional networks. Our participant pool comprises experts from $5$ universities and $2$ industrial labs, with $69\%$ being Ph.D. students or recent graduates and $31\%$ being associate researchers in industrial or academic labs. [^3] Our experts are affiliated with University of Washington, Carnegie Mellon University, Harvard University, University of Southern California, Indian Institute of Science, Allen Institute for Artificial Intelligence (Ai2) and Adobe, Inc. Unlike prior studies where participants assess LLM-generated research documents without suspecting plagiarism, our participants are instructed to actively search for potential overlaps with prior work.
@@ 5 Results
Our code and detailed findings, including all evaluated proposals and author verifications, are available at https://github.com/tarun360/AI-Papers-Plagiarism/.
@@ 5.1 Expert Evaluation Results
Our expert evaluation reveals plagiarism in LLM-generated research documents, as shown in Table 2. Similarity scores for verified claims highlight substantial content misappropriation at various levels of severity. Since our evaluation is limited by our participants’ time constraints and the laborious manual effort required for thorough plagiarism checks, our results likely represent a lower bound on actual plagiarism rates. Interestingly, of the four exemplars presented in Si et al. (2024), one received a similarity score of $5$ and another received a score of $4$, while among the ten exemplars in Lu et al. (2024a), two received scores of $5$ and one received score $4$—all of these ratings are cross-verified by the source papers’ authors. It is crucial to note the documents that receive similarity scores of $5$ in our evaluation passed through SSAG plagiarism detection checks, indicating serious limitations of such systems (elaborated in §5.3).
@@
5.2 A Case Study
As shown in Figure 2, the proposed methodology exhibits a clear one-to-one mapping with the original paper. Each component of the LLM-generated proposal corresponds to specific sections in the source paper, albeit with skillfully reworded descriptions. The proposal proposes the same technical approach while using different terminology (e.g., “resonance graph” instead of “weighted adjacency matrix”) and restructuring the presentation. This may be interpreted as adversarial behavior, where the LLM has learned to disguise existing work as novel research through careful rewording. Notably, expert reviewers in Si et al. (2024)’s study do not identify this plagiarism, likely because their evaluation focuses on assessing novelty and feasibility rather than actively searching for potential sources of plagiarism. In contrast, our study’s participants, operating under different situational logic (Popper, 2013; Hoover, 2016) that presumes potential plagiarism, are able to identify the original paper. This case-study, along with results of expert evaluation presented in §5.1, showcases the limitations of previous human evaluations of LLM generated research ideas—without a skeptical eye for plagiarism, experts may be fooled into considering LLM generated research ideas as novel. Additional case studies examining plagiarism and similarity in AI-generated research from (Lu et al., 2024a) and (Yamada et al., 2025) are available in §A and §B respectively.
5.3 Performance of Plagiarism Detectors
@@
The SSAG approach, a common plagiarism detection method in research document generation systems (Si et al., 2024; Lu et al., 2024a; Li et al., 2024a), attains moderate performance ($51.3\%$ for Claude, $68.5\%$ for GPT-4o). The considerable gap between SSAG and oracle access performance indicates that retrieving relevant papers, not determining similarity, is the bottleneck.
SSAG’s inadequate performance is particularly concerning for research agent systems relying on it (Si et al., 2024; Lu et al., 2024a; Li et al., 2024a), especially given our test scenario represents a much easier challenge—our proposals are deliberately plagiarized from single papers, while research agent systems are designed to produce allegedly novel content. (An illustrative example of successful plagiarism detection using the SSAG from our synthetic dataset is provided in Appendix C.)
@@ 6 Discussion
Comparison with Human-Written Papers.
In principle, it would be useful to conduct a similar analysis of human-written papers to estimate rates of plagiarism in them. However, given the labor-intensive nature of finding plagiarism in LLM-generated papers, our main study focuses on analyzing LLM-generated content alone. As a proxy to the expert-led study on human-written papers, we conduct an automated analysis to extract signs of plagiarism from peer reviews of human-written papers. We use existing peer reviews from the PeerRead dataset (Kang et al., 2018) for human-written papers from NeurIPS 2017 ($499$ papers), ACL 2017 ($123$ papers), ICLR 2017 ($349$ papers), and CoNLL 2016 ($19$ papers). For NeurIPS 2017, we sampled $499$ papers from the larger set due to computational constraints. We use Claude 3.7 Sonnet to extract potential concerns of plagiarism mentioned in peer reviews, using the prompt shown in Table 11. We emphasize that the LLM is used only to extract information about reviewer claims, not to make plagiarism judgments. As shown in Table 4, the estimated plagiarism rate in human-authored papers is significantly lower than the $24\%$ we found in LLM-generated papers. This stark difference reinforces our finding that LLM-generated research exhibits substantially higher rates of content misappropriation than traditional human-authored work.
Conference Score $4$ Score $5$ Plagiarism
rate (scores $4$+)
ACL 2017 $0.8\%$ $0\%$ $0.8\%$
ICLR 2017 $4.0\%$ $2.3\%$ $6.3\%$
CoNLL 2016 $5.3\%$ $0\%$ $5.3\%$
NeurIPS 2017 $1.8\%$ $0\%$ $1.8\%$
Table 4: We estimate plagiarism rates in human-written papers by automatically extracting signs of plagiarism from peer reviews. We find that the overall plagiarism rate is significantly lower than the $24\%$ found in LLM-generated proposals.
Figure 3: PCA projection of concatenated title and abstract embeddings for human-written papers and LLM-generated proposals on the topic “Novel AI-assisted formal proof generation methods”. LLM-generated proposals occupy a more confined region, indicating less diversity in outputs.
Limited Diversity in Proposals.
Prior research notes limited diversity in LLM-generated research proposals (Si et al., 2024), which our analysis confirms. For each research topic in Table 5, we analyze concatenated titles and abstracts of $100$ human-written papers and $100$ LLM-generated proposals. [^5] For these experiments, we modify Si et al. (2024)’s proposal generation to output title and abstract format (instead of the original structure described in §3) to enable direct comparison with human-written papers. Using the all-MiniLM-L6-v2 sentence transformer (Reimers and Gurevych, 2020), we generate embeddings for these concatenated titles and abstracts. Calculating cluster spread via mean squared distance from centroids, we find that despite aggressive deduplication strategies (Si et al., 2024), proposals generated by both LLMs are more tightly clustered—the ratio of LLM to human cluster spreads is $0.81$ for proposals generated by Claude 3.5 Sonnet and $0.72$ for proposals generated by GPT-4o across topics. [^6] Due to deduplication and filtering in (Si et al., 2024), some topics have $<100$ proposals, resulting in $2370$ samples for Claude 3.5 Sonnet and $2019$ samples for GPT-4o rather than the maximum of $2400$ ($200$ samples $\times$ $12$ topics). Figure 3 illustrates this pattern, showing that LLM proposals occupy a more confined region in the embedding space. While this analysis is limited since titles and abstracts only reflect broad research directions rather than specific ideas, the clear clustering pattern suggests current LLM systems explore a narrower range of research directions compared to humans, undermining their utility for scientific innovation.
Proposals Are Easily Distinguishable.
Beyond being tightly clustered, LLM-generated proposals also appear to follow distinctive patterns in their research directions. In a preliminary analysis using $100$ human-written papers and $100$ LLM-generated proposals per topic from Table 5, we train separate binary classifiers to distinguish between human-written papers and LLM-generated proposals from Claude 3.5 Sonnet, and between human-written papers and proposals from GPT-4o.5 We use a simple logistic regression model with all-MiniLM-L6-v2 sentence transformer (Reimers and Gurevych, 2020) features on a $60$:$40$ train-test split. The classifiers achieve $93.9\%$ accuracy (Claude 3.5 Sonnet vs. human) and $92.6\%$ accuracy (GPT-4o vs. human).6 While this basic classification experiment has limitations as titles and abstracts only capture broad research directions, it provides initial evidence that LLM-generated research follows predictable patterns. However, developing reliable detection methods will require significant further research and is beyond the scope of this work.
Citing Relevant Work Is Challenging.
@@ 8 Limitations
Our expert evaluation design has some limitations. First, we ask our human participants to presume adversarial plagiarism and actively search for it. This may introduce confirmation bias, leading them to give higher similarity scores in order to complete their task. We indeed find $4$ instances where source-papers authors over-turned the score downwards with differences of $2$ or more. However, our (author-verified) scores already account for these adjustments, and our findings remain independently verifiable through our open-sourced results. Second, we provide our human-participants with $5$ research proposals and ask them to choose any $3$. While this gives flexibility to our participants, since even in a single topic there can be sub-topics that participants may not be familiar with, it may introduce sampling bias.